OpenStudy (anonymous):

a ball is thrown vertically downward with a speed of 3.0m/s.find its speed when it has travelled a distance of 5m.

OpenStudy (y2o2):

you have to state the gravitational acceleration.

OpenStudy (anonymous):

i think the gravitational acceleration is 9.81m/s^2. a question is given like that without state the gravitational acceleration.

OpenStudy (y2o2):

So you'll use the third Equation of motion: v² = u² + 2as so v² = 3² + 2*9.81*5 =107.1 v = sqrt(v²) = 10.349 m/s

OpenStudy (anonymous):

tq very much..i'm appreciate it..

OpenStudy (y2o2):

I'm glade it helped :)